OpenStudy (anonymous):

Select all of the following statements that are true. 6 + 3 = 9 and 4 · 4 = 16 If 6 · 3 = 18, then 4 + 8 = 20 5 · 3 = 15 or 7 + 5 = 20 If 6 > 10, then 8 · 3 = 24

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

ok, is statement 1 true or not?

OpenStudy (anonymous):

1st is, 2nd isn't, 3rd also isnt... but im not sure about the last one

OpenStudy (solomonzelman):

I am not sure if you are correct about the 3rd one. The way I see the third statement is that (since it says "or") you need one of the 2 to be true for the statement to be valid. either \(\large\color{slate}{ 5\times 3=15}\) or \(\large\color{slate}{ 5+5=20}\)

OpenStudy (solomonzelman):

Is any of the things in the third statement true?

OpenStudy (anonymous):

yep. the 5x3=15 is

OpenStudy (anonymous):

so its true?

OpenStudy (solomonzelman):

Yes (apparently). it is "or" and thus, only one thing has to be true. (if it was "and" then both things would have to be true, and then your statement would have been false).

OpenStudy (anonymous):

so that means the 4th one is also true?

OpenStudy (solomonzelman):

in statement 4 there is no "or"

OpenStudy (solomonzelman):

Statement 4. If \(\large\color{slate}{ 6>10}\), then \(\large\color{slate}{8 \cdot 3=24}\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ 6>10}\) ? yes or no

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

So statement 4 is ?

OpenStudy (anonymous):

false?

OpenStudy (solomonzelman):

Yup

OpenStudy (anonymous):

but only 1 and 3 are incorrect

OpenStudy (solomonzelman):

only 1 is a true statement.

OpenStudy (anonymous):

ohh ok thnx

OpenStudy (solomonzelman):

yes...

OpenStudy (solomonzelman):

yw